OpenStudy (beccab003):

Simplifying radicals help! Medal to best answer!! Question will be posted below.

OpenStudy (beccab003):

\[\sqrt[3]{4/9}\]

OpenStudy (beccab003):

It needs to be simplified and I'm not sure how to do that with a fraction.

OpenStudy (beccab003):

@Directrix could you help maybe?

OpenStudy (campbell_st):

rewrite them in index form \[\sqrt[3]{\frac{4}{9}} = \sqrt[3]{(\frac{2}{3})^2} = \sqrt[3]{2^2 3^{-2}}\] now apply the index rule for the cube roots... \[\sqrt[a]{x} = x^{\frac{1}{a}}\]

OpenStudy (beccab003):

Thanks campbell! So now I do: \[\sqrt[3]{2^2*1/3 * 3^2*1/3}\]

OpenStudy (beccab003):

@campbell_st

OpenStudy (campbell_st):

no its \[(2^2)^{\frac{1}{3}} \times (3^{-2})^{\frac{1}{3}}\] just simplify this

OpenStudy (beccab003):

but it doesn't come out as a whole number?

OpenStudy (beccab003):

I am just supposed to simplify.

OpenStudy (campbell_st):

you need the index law for power of a power.... which requires you to multiply the powers \[(x^a)^b = x^{a \times b}\] you need to use this law to finish this question

OpenStudy (beccab003):

so like this? \[2^{2/3}*3^{-2/3}\]

OpenStudy (beccab003):

@campbell_st

OpenStudy (campbell_st):

that's correct.... it can be written with out the multiplication.

OpenStudy (beccab003):

I need the final answer to be written as a fraction and with a square root. Can you help with that? thanks.

OpenStudy (beccab003):

@campbell_st

OpenStudy (campbell_st):

if its as a fraction with a cube root its simply \[\sqrt[3]{\frac{2^2}{3^2}}\]

OpenStudy (beccab003):

That is not one of my options....here they are: \[\frac{ \sqrt[3]{12} }{ 3 }\] \[\frac{ \sqrt[3]{324} }{ 9 }\] \[\frac{ \sqrt[3]{5} }{ 3 }\] \[\frac{ \sqrt[3]{36} }{ 9 }\]

OpenStudy (beccab003):

@campbell_st

OpenStudy (beccab003):

@Directrix I really need some assistance. Thanks!

OpenStudy (beccab003):

@mathstudent55 @dan815 @Mimi_x3 @Abhisar @TheSmartOne @CGGURUMANJUNATH Can someone help me please?

Directrix (directrix):

Where is the question? @BeccaB003

OpenStudy (beccab003):

Here it is: Simplify: \[\sqrt[3]{\frac{ 4 }{ 9 }}\]

OpenStudy (beccab003):

the answer options are the fifth one up that I already posted.

Directrix (directrix):